OpenStudy (alexh107):

Find the solutions to a^3+5a=15a^2

OpenStudy (itz_sid):

\[a^3+5a=15a^2\] So you start by bringing everything to one side.\[a^3-15a^2+5a\] You can start by factoring out an "a" from the whole thing. So you get...\[a(a^2-15a+5)\] which means a=0. And then you factor \[(a^2-15a+5a)\] and set it equal to zero to find the rest of your solutions.

OpenStudy (alexh107):

I managed to get that far and know 0 is one of the solutions, but I don't know how to get the other two.

OpenStudy (itz_sid):

Oh okay, I see. You cannot factor that using the guess and check method, so you are going to have to use the Quadratic Equation Formula: x = −b ± √(b2 − 4ac) 2a.

OpenStudy (alexh107):

Ah okay, I was just thinking that. I'll work that out now.

OpenStudy (alexh107):

Okay, I found the answer as 0, \[(15+-\sqrt{205})/2\]